OpenStudy (itsonlycdeee):

I'm not sure if the definite integral of these two parts is 0+0=0 or -1/4 + 1/4 = 0. Question below:

OpenStudy (itsonlycdeee):

\[\int\limits_{-\infty}^{\infty} x^3e^-x^4 dx\]

OpenStudy (itsonlycdeee):

\[\int\limits_{-\infty}^{0} x^3e^-x^4 dx = \lim_{a \rightarrow -\infty} \int\limits_{a}^{0}x^3e^-x^4= x^3e^-x^4 \] | a --> 0

OpenStudy (itsonlycdeee):

\[\int\limits_{0}^{\infty} x^3e^-x^4 dx = \lim_{b \rightarrow \infty} \int\limits_{0}^{b} x^3e^-x^4 = x^3e^-x^4 | 0\rightarrow b\]

OpenStudy (anonymous):

is it \[\int x^3e^{x^4}dx\]?

OpenStudy (anonymous):

\[\large \int x^3e^{-x^4}dx\]

OpenStudy (anonymous):

if so, it is an odd function

OpenStudy (itsonlycdeee):

@satellite73 Yes, its the second integral you posted with a negative in front of x^4.

ganeshie8 (ganeshie8):

yes, the definite integral equals 0... but u seem to be asking a different question... can u elaborate a bit ha ? :)

OpenStudy (itsonlycdeee):

& thank you.

OpenStudy (itsonlycdeee):

@ganeshie8 * Oh, okay aha. Well my question was the way I got the definite integral as 0. Like the two ways I posted it. Are each integral, when done separately 0 + 0 which equals 0? or what my classmate got, which was -1/4 + 1/4 = 0?

ganeshie8 (ganeshie8):

\( \large \mathbb { \int\limits_{-\infty}^{0} x^3e^{-x^4} dx = \lim \limits_{a \rightarrow -\infty} \int\limits_{a}^{0}x^3e^{-x^4} dx}\)

ganeshie8 (ganeshie8):

First, evaluate the integral and see wat u get, u need to use substitution to evaluate the integral

ganeshie8 (ganeshie8):

dont wry about the limit yet... first evaluate the integral

OpenStudy (itsonlycdeee):

Oh, I used substitution & got. u = -x^4 du = -4x^3 

OpenStudy (itsonlycdeee):

\[\frac{ -1 }{ 4 } \int\limits_{}^{} e ^{-u} du\]

ganeshie8 (ganeshie8):

u = -x^4 du = -4x^3 dx -du/4 = x^3 dx

ganeshie8 (ganeshie8):

^^

ganeshie8 (ganeshie8):

yes ! that looks good... keep going...

OpenStudy (itsonlycdeee):

Yes that, sorry I'm reading my paper too fast. lol

ganeshie8 (ganeshie8):

:) u need to setup the bounds also ok ?

OpenStudy (itsonlycdeee):

Ok, so [0, infinity] its \[\frac{ -1 }{ 4 }e ^{-x^4}(-4x ^{3}) +C \] ?

ganeshie8 (ganeshie8):

\(\large \mathbb { \int\limits_{-\infty}^{0} x^3e^{-x^4} dx = \lim \limits_{a \rightarrow -\infty} \int\limits_{a}^{0}x^3e^{-x^4} dx} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \lim \limits_{a \rightarrow -\infty} \int\limits_{a^4}^{0}e^{-u} (\frac{du}{4})} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \lim \limits_{a \rightarrow -\infty} \frac{1}{4} \int\limits_{a^4}^{0}e^{-u} du} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \lim \limits_{a \rightarrow -\infty} \frac{1}{4} -e^{-u}\Big|_{a^4}^{0}} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \lim \limits_{a \rightarrow -\infty} \frac{1}{4} [-e^{-0} + e^{-a^4} ]} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \lim \limits_{a \rightarrow -\infty} \frac{1}{4} [-1 + e^{-a^4} ]} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \frac{1}{4} [-1 + e^{-\infty} ]} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \frac{1}{4} [-1 + 0]} \) \(\large \mathbb { ~~~~~~~~~~~~~~~~~~~ = \frac{-1}{4}} \)

ganeshie8 (ganeshie8):

^^complete solution above.. see if that makes some sense...

ganeshie8 (ganeshie8):

similarly the integral other side equals 1/4 both adding gives u -1/4 + 1/4 = 0 as it should be for an odd function

OpenStudy (itsonlycdeee):

Ok, thank you! I'm just confused as to why the bound is [a^4, 0], instead of [a, 0]

ganeshie8 (ganeshie8):

good question :) we substituted u = x^4 so, as x -> a, u -> a^4 as x -> 0, u -> 0

OpenStudy (itsonlycdeee):

Ohhhh, okay. I get it now. Thank you very much!

ganeshie8 (ganeshie8):

np :) try to do the other integral also... and see if u get 1/4.. good luck !